OpenStudy (anonymous):

let f(x,y) = square-root(1+4x+y^2) and let P be the point (1,2) a) at P what is the direction of maximal increase for the function f? give your answer as a init vector.

ganeshie8 (ganeshie8):

what do you know about gradient vector ?

OpenStudy (anonymous):

You want to maximize \(df\).

OpenStudy (anonymous):

I know that is \[f_x i +f_y j + f_z z\]

OpenStudy (anonymous):

Get \(df\) using the chain rule.

OpenStudy (anonymous):

the integral of f(x,y)?

OpenStudy (anonymous):

No, the derivative

OpenStudy (anonymous):

derivative to respect to y or x?

ganeshie8 (ganeshie8):

griadent points in the maximum increasing direction of the function so you just need to find the gradient at given point

ganeshie8 (ganeshie8):

the given function is in two variables, so gradient would be just \[\langle f_x, ~f_y\rangle \]

OpenStudy (anonymous):

do i have to sub in the point P?

ganeshie8 (ganeshie8):

yes you want gradient at the given point

ganeshie8 (ganeshie8):

gradient is a vector field so you will be having one vector at each point

OpenStudy (anonymous):

\[gradf = 2/(\sqrt{y^2+4y+1}+y/\sqrt{4x+y^2+1}\]

ganeshie8 (ganeshie8):

thats not a vector ^

OpenStudy (anonymous):

\[grad f = \frac{ 2 }{ \sqrt{y^2+4y+1}} i + \frac{ y }{ \sqrt{4x+y^2+1} } j\]

ganeshie8 (ganeshie8):

looks better :)

OpenStudy (anonymous):

did i get the gradient right?

ganeshie8 (ganeshie8):

looks you have few typoes

OpenStudy (anonymous):

where?

ganeshie8 (ganeshie8):

in the denominator expression

ganeshie8 (ganeshie8):

\[grad f = \frac{ 2 }{ \sqrt{1+4x+y^2}} i + \frac{ y }{ \sqrt{1+4x+y^2} } j \] right ?

OpenStudy (anonymous):

oh ok I see where I made a mistake

OpenStudy (anonymous):

so I have to sub in the point (1,2)

ganeshie8 (ganeshie8):

yep that gives you gradient vector at that particular point

OpenStudy (anonymous):

\[\frac{ 2 }{ \sqrt{9} } i + \frac{ 2 }{ \sqrt{9} } j \]

ganeshie8 (ganeshie8):

looks good ! read the question again, they want unit vector so divide by the magnitude

OpenStudy (anonymous):

divide by the magnitude of the gradf?

ganeshie8 (ganeshie8):

hats the unit vector in the direction of \[\frac{ 2 }{ \sqrt{9} } i + \frac{ 2 }{ \sqrt{9} } j\] ?

OpenStudy (anonymous):

\[ df = \frac{\partial f}{\partial x}dx + \frac{\partial f}{\partial y}dy = \left\langle \frac{\partial f}{\partial x},\frac{\partial f}{\partial y}\right\rangle \cdot d\langle x,y\rangle \]We can always let \( d\langle x,y\rangle\) be parallel to \( \left\langle \frac{\partial f}{\partial x},\frac{\partial f}{\partial y}\right\rangle \) to maximize \(df\).

ganeshie8 (ganeshie8):

yeah we are maximizing \[\nabla f \cdot \dfrac{d\vec{r}}{dt} \]

ganeshie8 (ganeshie8):

\[\frac{ 2 }{ \sqrt{9} } i + \frac{ 2 }{ \sqrt{9} } j\] Notice this points in the same direction as \[1i + 1j\]

OpenStudy (anonymous):

so that could be the unit vector?

OpenStudy (anonymous):

i + j

ganeshie8 (ganeshie8):

i+j is NOT an unit vector

ganeshie8 (ganeshie8):

the magnitude of unit vector has to be 1

ganeshie8 (ganeshie8):

whats the magnitude of vector i+j ?

OpenStudy (anonymous):

2

ganeshie8 (ganeshie8):

magnitude of ai+bj is the distance from origin to point (a, b)

OpenStudy (anonymous):

squareroot (2)

ganeshie8 (ganeshie8):

so whats the unit vector in the direction of i+j ?

OpenStudy (anonymous):

squareroot(1^2 +1^2)

ganeshie8 (ganeshie8):

the unit vector in the direction of i+j would be \[\dfrac{1}{\sqrt{1^2+1^2}}i + \dfrac{1}{\sqrt{1^2+1^2}} j\]

OpenStudy (anonymous):